# Helicity controls the direction of fluxes in rotating turbulence

### v1

@@ Appendix D Heterochiral interactions: governing equations
and that $\langle v_{-\bm{p}}g_{\bm{p},y}\rangle=0$ to derive Eq. (71b). In steady state, Eqs. (66)-(68) lead to
@@
$\partial_{p_{y}}\Big(p^{4}\Re(\langle vv\rangle_{\bm{p}\bm{k}})\Big)$ $=-p^{4}\frac{\sqrt{2}\epsilon\chi_{\bm{p},y}}{p_{x}{U^{\prime}}}$ (71a)
$\partial_{p_{y}}\Big(p^{2}\langle\eta v\rangle_{\bm{p}\bm{k}}\Big)$ $=ip^{2}\frac{p_{z}}{p_{x}}\langle vv\rangle_{\bm{p}\bm{k}}$ (71b)
@@
$\partial_{p_{y}}\Re(\langle\eta\eta\rangle_{\bm{p}\bm{k}})$ $=2\frac{p_{z}}{p_{x}}\Im\langle\eta v\rangle_{\bm{p}\bm{k}}-\frac{\sqrt{2}\epsilon p^{2}\chi_{\bm{p},y}}{p_{x}{U^{\prime}}},$ (71c)
@@
after simple algebraic manipulations, and noting that $\Im(\langle vv\rangle_{\bm{p}\bm{k}})$ and $\Im(\langle\eta\eta\rangle_{\bm{p}\bm{k}})$ are $O(k_{y}/p)$.
Note that, owing to incompressibility, $u_{\bm{p}}$ and $w_{\bm{p}}$ are connected to $v_{\bm{p}}$ and $\eta_{\bm{p}}$ via
$u_{\bm{p}}=-i\frac{p_{z}}{p_{0}^{2}}\eta_{\bm{p}}-\frac{p_{x}p_{y}}{p_{0}^{2}}v_{\bm{p}},~~~~~~~~~~w_{\bm{p}}=i\frac{p_{x}}{p_{0}^{2}}\eta_{\bm{p}}-\frac{p_{y}p_{z}}{p_{0}^{2}}v_{\bm{p}},$ (72)
@@
from which we can express the Reynolds stress $\langle uv\rangle_{\bm{p}\bm{k}}$ as a function of correlators $\langle vv\rangle_{\bm{p}\bm{k}}$ and $\langle\eta v\rangle_{\bm{p}\bm{k}}$:
$\langle uv\rangle_{\bm{p}\bm{k}}=-\frac{p_{x}p_{y}}{2(p_{x}^{2}+p_{z}^{2})}\langle vv\rangle_{\bm{p}\bm{k}}+\frac{p_{z}}{2(p_{x}^{2}+p_{z}^{2})}\Im\langle\eta v\rangle_{\bm{p}\bm{k}}.$ (73)
@@
One can further associate variables $(v_{\bm{p}},\eta_{\bm{p}})$ to the helical variables $(a_{\bm{p}}^{+},a_{\bm{p}}^{-})$, via basis (23). For the correlators, this results in applying the transform
$\langle vv\rangle_{\bm{p}\bm{k}}$ $=\frac{p_{0}^{2}}{2p^{2}}(\Phi_{\bm{p}\bm{k}}^{+}+\Phi_{\bm{p}\bm{k}}^{-})~~~+\Re(Z)(\Psi_{\bm{p}\bm{k}}^{+-}+\Psi_{\bm{p}\bm{k}}^{-+})+i\Im(Z)(\Psi_{\bm{p}\bm{k}}^{+-}-\Psi_{\bm{p}\bm{k}}^{-+})$ (74)
$\langle\eta v\rangle_{\bm{p}\bm{k}}$ $=p\frac{p_{0}^{2}}{2p^{2}}(\Phi_{\bm{p}\bm{k}}^{+}-\Phi_{\bm{p}\bm{k}}^{-})+p(Z\Psi_{\bm{p}\bm{k}}^{+-}-Z^{*}\Psi_{\bm{p}\bm{k}}^{-+})$ (75)
$\langle\eta\eta\rangle_{\bm{p}\bm{k}}$ $=\frac{p_{0}^{2}}{2}(\Phi_{\bm{p}\bm{k}}^{+}+\Phi_{\bm{p}\bm{k}}^{-})~~~-p^{2}\Re(Z)(\Psi_{\bm{p}\bm{k}}^{+-}+\Psi_{\bm{p}\bm{k}}^{-+})-ip^{2}\Im(Z)(\Psi_{\bm{p}\bm{k}}^{+-}-\Psi_{\bm{p}\bm{k}}^{-+})$ (76)
@@
to system (40)-(41), with $p_{0}^{2}=p_{x}^{2}+p_{z}^{2}$ and $Z=\frac{p_{z}^{2}p_{y}^{2}-p_{x}^{2}p^{2}}{2p^{2}p_{\perp}^{2}}+i\frac{2p_{x}p_{y}p_{z}}{2p_{\perp}^{2}p}$. With zero helicity injection, $\Psi_{\bm{p}\bm{k}}^{+-}=\Psi_{\bm{p}\bm{k}}^{-+*}$ and $\Phi_{\bm{p}\bm{k}}^{+}=\Phi_{\bm{p}\bm{k}}^{-}$, and the transformation is simplified. Equations (66)-(68) can then be directly obtained from Eqs. (40)-(41) through this transformation (we do not detail the computation here).
@@
Note that in the absence of cross-chirality correlations $\Psi_{\bm{p}\bm{k}}^{+-}=0$, and if both helicity signs are excited isotropically ($\Phi_{\bm{p}\bm{k}}^{+}=\Phi_{\bm{p}\bm{k}}^{-}$), the correlation between velocity and vorticity of different waves vanishes, $\langle\eta v\rangle_{\bm{p}\bm{k}}=0$. As a result, velocity and vorticity are enslaved in Sector H:
$\langle vv\rangle_{\bm{p}\bm{k}}=\frac{p_{0}^{2}}{p^{2}}\Phi_{\bm{p}\bm{k}}^{s}=\frac{\langle\eta\eta\rangle_{\bm{p}\bm{k}}}{p^{2}}.$ (77)
@@ Appendix E Flux coexistence due to rotation
dictates which heterochiral-wave interactions with the condensate are possible. Note that the rotation rate also enters in the resonant condition for the homochiral interactions, via filter $F_{k_{f}}$ (59), but here we fix $F_{k_{f}}=1$: we consider the regime where all homochiral-wave interactions contribute and do not start to decouple due to stringent resonances. This regime is analyzed in [gome2025waves].
@@
Our goal here is to derive the energy transfer from the fluctuations to the mean flow, and its leading-order behavior with variable $S$ and domain geometry. Due to the conservation of energy, this total transfer can be computed either by integrating the Reynolds stress transfer ${U^{\prime}}\langle uv\rangle_{\bm{p}\bm{k}}$, or by computing the difference between the total energy flux to small scales and the energy injection rate per mode $\epsilon_{p_{x},p_{z}}$ (assuming $F_{k_{f}}=1$):
$\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{p_{x},p_{z}}$ $\equiv\int dp_{y}\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\bm{p}\bm{k}}$ (81)
@@
The exact computation of the spectral energy flux $\Pi_{\bm{p}}$ in Layers II was carried out in the End Matter (see Eq. (18)). In §E.1, we will show that the $\bm{p}$-integrated energy flux originates from these layers, producing a dominant forward energy transfer in $O\left(S\frac{L_{x}}{l_{f}}\right)$. Then, we derive the exact Reynolds stress in Layers I, and compute subleading terms coming from both Layers I and II, which we will show are in $O(S)$ and $O(\frac{L_{x}}{l_{f}}\sqrt{\frac{l_{f}}{L_{z}}})$. These terms, in part due to discretization errors, will be included to better compare with our DNS setup.
@@ E.1 Leading-order energy transfer
$\Pi_{\bm{p}}=\frac{\Pi^{*}}{2}\Bigg[1+\frac{p^{*2}}{p^{2}}+\frac{p^{*2}}{p_{0}^{2}}\frac{p_{z}^{2}}{p_{x}^{2}}\Big(\arctan\Big(\frac{p_{y}}{p_{0}}\Big)-\arctan\Big(\frac{p_{y}^{*}}{p_{0}}\Big)\Big)^{2}\Bigg].$ (83)
The $\bm{p}$-dependent Reynolds transfer associated to Layers II is:
$\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\bm{p}\bm{k}}$ $=\sqrt{2}{U^{\prime}}\left(-\frac{p_{x}p_{y}}{p_{0}^{2}}\langle vv\rangle_{\bm{p}\bm{k}}+\frac{p_{z}}{p_{0}^{2}}\Im\langle\eta v\rangle_{\bm{p}\bm{k}}\right)$ (84)
@@
With zero helicity injection, and if all homochiral-wave interactions are resonant ($F_{k_{f}}=1$), the total energy balance between the small-scale flux $\Pi_{\bm{p}}$ (83) and the transfer to the mean flow from layers II is therefore:
$\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{p_{x},p_{z}}^{\rm II}\equiv\int dp_{y}\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\bm{p}\bm{k}}^{\rm II}$ $=\epsilon_{p_{x},p_{z}}-\Pi_{p_{x},K_{y},p_{z}},$ (86)
@@
We now extract the leading order functional form of the Reynolds stress in the variable $S=\frac{{U^{\prime}}}{\Omega}$. Here we consider that $S$ is sufficiently large ($S\gtrsim l_{f}/L_{y}$, following the computation in [gome2025waves]) so that all the waves excited at the forcing scale are coupled with the condensate via homochiral interactions, hence $F_{k_{f}}=1$. We also assume that
@@
The energy transfer between the waves and the 2D condensate in (105) decreases with decreasing rotation, via increasing $S={U^{\prime}}/\Omega$, due to the growing number of modes in sector A that break conservation of $H_{\bm{p}}^{\pm}$ and extract mean-flow energy. This energy extraction is larger as $L_{x}/l_{f}$ increases, due to vortex-tilting preferentially energizing modes with large $p_{z}/p_{x}\sim L_{x}/l_{f}$, via the term in $p_{z}/p_{x}$ in (87).
@@ E.2.1 Reynolds stress in Layers I
$\langle vv\rangle_{\bm{p}\bm{k}}$ $=\frac{\epsilon_{p_{x},p_{z}}}{2\sqrt{2}p_{x}{U^{\prime}}}\frac{p_{0}^{2}k_{f}^{2}}{p^{4}},$
$\langle\eta v\rangle_{\bm{p}\bm{k}}$ $=i\frac{\epsilon_{p_{x},p_{z}}}{2\sqrt{2}p_{x}{U^{\prime}}}p_{0}\frac{p_{z}}{p_{x}}\frac{k_{f}^{2}}{p^{2}}\Bigg[\arctan\Big(\frac{p_{y}}{p_{0}}\Big)+\arctan\Big(\frac{p_{y}^{f}}{p_{0}}\Big)\Bigg],~~~~~~~-p_{y}^{f}<p_{y}<-p_{y}^{*}$ (108)
@@
$\Phi_{\bm{p}\bm{k}}^{s}=\frac{\epsilon_{p_{x},p_{z}}}{2p_{x}{U^{\prime}}}\frac{k_{f}^{2}}{p^{*}p},$ (109)
using the continuity of the flux at $p_{y}^{*}$, $\Phi_{\bm{p}\bm{k}}^{s}=\frac{p^{*2}}{p_{0}^{2}}\langle vv\rangle_{\bm{p}\bm{k}}(-p_{y}^{*})$. Because Eq. (53) exhibits $p_{y}\to-p_{y}$ symmetry, the boundary condition at $p_{y}^{*}$ dictates the values of the correlators at $p_{y}^{*}$.
@@
$\langle vv\rangle_{\bm{p}\bm{k}}=\langle vv\rangle_{\bm{p}\bm{k}}(-p_{y}^{*})\frac{p^{*4}}{p^{4}}+\frac{\sqrt{2}\epsilon}{2|{U^{\prime}}p_{x}|p^{4}}\int_{p_{y}^{*}}^{p_{y}}q^{4}(1-\frac{q_{y}^{2}}{q^{2}})\chi_{\bm{q}}dq_{y}=\left\{\begin{aligned} &\frac{\epsilon_{p_{x},p_{z}}}{2\sqrt{2}|{U^{\prime}}p_{x}|}\frac{p_{0}^{2}k_{f}^{2}}{p^{4}},~~~p_{y}^{*}<p_{y}<p_{y}^{f}\\ &\frac{\epsilon_{p_{x},p_{z}}}{\sqrt{2}|{U^{\prime}}p_{x}|}\frac{p_{0}^{2}k_{f}^{2}}{p^{4}},~~~p_{y}>p_{y}^{f}\end{aligned}\right$. (110)
$\langle\eta v\rangle_{\bm{p}\bm{k}}=\langle\eta v\rangle_{\bm{p}\bm{k}}(-p_{y}^{*})\frac{p^{*2}}{p^{2}}+\frac{i}{p^{2}}\int_{p_{y}^{*}}^{p_{y}}\frac{p_{z}}{p_{x}}q^{2}v_{q}^{2}dq_{y}=\left\{\begin{aligned} &i\frac{\epsilon_{p_{x},p_{z}}~p_{0}}{2\sqrt{2}{U^{\prime}}p_{x}}\frac{p_{z}}{p_{x}}\frac{k_{f}^{2}}{p^{2}}\Bigg[\arctan(\frac{p_{y}}{p_{0}})+\arctan(\frac{p_{y}^{f}}{p_{0}})-2\arctan(\frac{p_{y}^{*}}{p_{0}})\Bigg],\\ &~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~p_{y}^{*}<p_{y}<p_{y}^{f}\\ &i\frac{\epsilon_{p_{x},p_{z}}p_{0}}{\sqrt{2}{U^{\prime}}p_{x}p_{y}^{f}}\frac{p_{z}}{p_{x}}\frac{k_{f}^{2}}{p^{2}}\Bigg[\arctan(\frac{p_{y}}{p_{0}})-\arctan(\frac{p_{y}^{*}}{p_{0}})\Bigg],\\ &~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~p_{y}>p_{y}^{f}\end{aligned}\right$.
@@
$\langle uv\rangle_{\bm{p}\bm{k}}=\left\{\begin{aligned} &\frac{\epsilon_{p_{x},p_{z}}k_{f}^{2}}{\sqrt{2}U^{\prime}}\Bigg[\frac{p_{y}}{p^{4}}-\frac{p_{z}^{2}}{p_{x}^{2}}\frac{1}{p_{0}}\frac{1}{p^{2}}\Big(\arctan\Big(\frac{p_{y}}{p_{0}}\Big)+\arctan\Big(\frac{p_{y}^{f}}{p_{0}}\Big)\Big)\Bigg],~~~~~-p_{y}^{f}<p_{y}<-p_{y}^{*}\\ &\frac{\epsilon_{p_{x},p_{z}}k_{f}^{2}}{\sqrt{2}U^{\prime}}\Bigg[\frac{p_{y}}{p^{4}}-\frac{p_{z}^{2}}{p_{x}^{2}}\frac{1}{p_{0}}\frac{1}{p^{2}}\Big(\arctan(\frac{p_{y}}{p_{0}})+\arctan(\frac{p_{y}^{f}}{p_{0}})-2\arctan(\frac{p_{y}^{*}}{p_{0}}\Big)\Bigg],~~~p_{y}^{*}<p_{y}<p_{y}^{f}\\ &\frac{\sqrt{2}\epsilon_{p_{x},p_{z}}k_{f}^{2}}{U^{\prime}}\Bigg[\frac{p_{y}}{p^{4}}-\frac{p_{z}^{2}}{p_{x}^{2}}\frac{1}{p_{0}}\frac{1}{p^{2}}\Big(\arctan(\frac{p_{y}}{p_{0}})-\arctan(\frac{p_{y}^{*}}{p_{0}}\Big)\Bigg],~~~~~~~p_{y}>p_{y}^{f}\end{aligned}\right$. (111)
In Fig. 1e in the main text, we visualize ${U^{\prime}}\langle uv\rangle_{\bm{p}\bm{k}}$ in the $(p_{y},p_{z})$ plane by showing Eq. (111) in layers I, completing the Reynolds stress in layers II ( Eq. (85)). In the figure, we fix $p_{x}=8\pi/L_{x}$ and $S=1$.

### v2

@@ Appendix D Heterochiral interactions: governing equations
and that $\langle v_{-\boldsymbol{p}}g_{\boldsymbol{p},y}\rangle=0$ to derive Eq. (71b). In steady state, Eqs. (66)-(68) lead to
@@
$\partial_{p_{y}}\Big(p^{4}\Re(\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}})\Big)$ $=-p^{4}\frac{\sqrt{2}\epsilon\chi_{\boldsymbol{p},y}}{p_{x}{U^{\prime}}}$ (71a)
$\partial_{p_{y}}\Big(p^{2}\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}\Big)$ $=ip^{2}\frac{p_{z}}{p_{x}}\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ (71b)
@@
$\partial_{p_{y}}\Re(\langle\eta\eta\rangle_{\boldsymbol{p}\boldsymbol{k}})$ $=2\frac{p_{z}}{p_{x}}\Im\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}-\frac{\sqrt{2}\epsilon p^{2}\chi_{\boldsymbol{p},y}}{p_{x}{U^{\prime}}},$ (71c)
@@
after simple algebraic manipulations, and noting that $\Im(\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}})$ and $\Im(\langle\eta\eta\rangle_{\boldsymbol{p}\boldsymbol{k}})$ are $O(k_{y}/p)$.
Note that, owing to incompressibility, $u_{\boldsymbol{p}}$ and $w_{\boldsymbol{p}}$ are connected to $v_{\boldsymbol{p}}$ and $\eta_{\boldsymbol{p}}$ via
$u_{\boldsymbol{p}}=-i\frac{p_{z}}{p_{0}^{2}}\eta_{\boldsymbol{p}}-\frac{p_{x}p_{y}}{p_{0}^{2}}v_{\boldsymbol{p}},~~~~~~~~~~w_{\boldsymbol{p}}=i\frac{p_{x}}{p_{0}^{2}}\eta_{\boldsymbol{p}}-\frac{p_{y}p_{z}}{p_{0}^{2}}v_{\boldsymbol{p}},$ (72)
@@
from which we can express the Reynolds stress $\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ as a function of correlators $\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ and $\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}$:
$\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}=-\frac{p_{x}p_{y}}{2(p_{x}^{2}+p_{z}^{2})}\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}+\frac{p_{z}}{2(p_{x}^{2}+p_{z}^{2})}\Im\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}.$ (73)
@@
One can further associate variables $(v_{\boldsymbol{p}},\eta_{\boldsymbol{p}})$ to the helical variables $(a_{\boldsymbol{p}}^{+},a_{\boldsymbol{p}}^{-})$, via basis (23). For the correlators, this results in applying the transform
$\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ $=\frac{p_{0}^{2}}{2p^{2}}(\Phi_{\boldsymbol{p}\boldsymbol{k}}^{+}+\Phi_{\boldsymbol{p}\boldsymbol{k}}^{-})~~~+\Re(Z)(\Psi_{\boldsymbol{p}\boldsymbol{k}}^{+-}+\Psi_{\boldsymbol{p}\boldsymbol{k}}^{-+})+i\Im(Z)(\Psi_{\boldsymbol{p}\boldsymbol{k}}^{+-}-\Psi_{\boldsymbol{p}\boldsymbol{k}}^{-+})$ (74)
$\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}$ $=p\frac{p_{0}^{2}}{2p^{2}}(\Phi_{\boldsymbol{p}\boldsymbol{k}}^{+}-\Phi_{\boldsymbol{p}\boldsymbol{k}}^{-})+p(Z\Psi_{\boldsymbol{p}\boldsymbol{k}}^{+-}-Z^{*}\Psi_{\boldsymbol{p}\boldsymbol{k}}^{-+})$ (75)
$\langle\eta\eta\rangle_{\boldsymbol{p}\boldsymbol{k}}$ $=\frac{p_{0}^{2}}{2}(\Phi_{\boldsymbol{p}\boldsymbol{k}}^{+}+\Phi_{\boldsymbol{p}\boldsymbol{k}}^{-})~~~-p^{2}\Re(Z)(\Psi_{\boldsymbol{p}\boldsymbol{k}}^{+-}+\Psi_{\boldsymbol{p}\boldsymbol{k}}^{-+})-ip^{2}\Im(Z)(\Psi_{\boldsymbol{p}\boldsymbol{k}}^{+-}-\Psi_{\boldsymbol{p}\boldsymbol{k}}^{-+})$ (76)
@@
to system (40)-(41), with $p_{0}^{2}=p_{x}^{2}+p_{z}^{2}$ and $Z=\frac{p_{z}^{2}p_{y}^{2}-p_{x}^{2}p^{2}}{2p^{2}p_{\perp}^{2}}+i\frac{2p_{x}p_{y}p_{z}}{2p_{\perp}^{2}p}$. With zero helicity injection, $\Psi_{\boldsymbol{p}\boldsymbol{k}}^{+-}=\Psi_{\boldsymbol{p}\boldsymbol{k}}^{-+*}$ and $\Phi_{\boldsymbol{p}\boldsymbol{k}}^{+}=\Phi_{\boldsymbol{p}\boldsymbol{k}}^{-}$, and the transformation is simplified. Equations (66)-(68) can then be directly obtained from Eqs. (40)-(41) through this transformation (we do not detail the computation here).
@@
Note that in the absence of cross-chirality correlations $\Psi_{\boldsymbol{p}\boldsymbol{k}}^{+-}=0$, and if both helicity signs are excited isotropically ($\Phi_{\boldsymbol{p}\boldsymbol{k}}^{+}=\Phi_{\boldsymbol{p}\boldsymbol{k}}^{-}$), the correlation between velocity and vorticity of different waves vanishes, $\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}=0$. As a result, velocity and vorticity are enslaved in Sector H:
$\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}=\frac{p_{0}^{2}}{p^{2}}\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s}=\frac{\langle\eta\eta\rangle_{\boldsymbol{p}\boldsymbol{k}}}{p^{2}}.$ (77)
@@ Appendix E Flux coexistence due to rotation
dictates which heterochiral-wave interactions with the condensate are possible. Note that the rotation rate also enters in the resonant condition for the homochiral interactions, via filter $F_{k_{f}}$ (59), but here we fix $F_{k_{f}}=1$: we consider the regime where all homochiral-wave interactions contribute and do not start to decouple due to stringent resonances. This latter regime is analyzed in [26].
@@
Our goal here is to derive the energy transfer from the fluctuations to the mean flow, and its leading-order behavior with variable $S$ and domain geometry. Due to the conservation of energy, this total transfer can be computed either by integrating the Reynolds stress transfer ${U^{\prime}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}$, or by computing the difference between the total energy flux to small scales and the energy injection rate per mode $\epsilon_{p_{x},p_{z}}$ (assuming $F_{k_{f}}=1$):
$\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{p_{x},p_{z}}$ $\equiv\int dp_{y}\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ (81)
@@
The exact computation of the spectral energy flux $\Pi_{\boldsymbol{p}}$ in Layers II was carried out in the End Matter (see Eq. (18)). In §E.1, we will show that the $\boldsymbol{p}$-integrated energy flux originates from these layers, producing a dominant forward energy transfer in $O\left(S\frac{L_{x}}{l_{f}}\right)$. Then, we derive the exact Reynolds stress in Layers I, and compute subleading terms coming from both Layers I and II, which we will show are in $O(S)$ and $O(\frac{L_{x}}{l_{f}}\sqrt{\frac{l_{f}}{L_{z}}})$. These terms, in part due to discretization errors, will be included to better compare with our DNS setup.
@@ E.1 Leading-order energy transfer
$\Pi_{\boldsymbol{p}}=\frac{\Pi^{*}}{2}\Bigg[1+\frac{p^{*2}}{p^{2}}+\frac{p^{*2}}{p_{0}^{2}}\frac{p_{z}^{2}}{p_{x}^{2}}\Big(\arctan\Big(\frac{p_{y}}{p_{0}}\Big)-\arctan\Big(\frac{p_{y}^{*}}{p_{0}}\Big)\Big)^{2}\Bigg].$ (83)
The $\boldsymbol{p}$-dependent Reynolds transfer associated to Layers II is:
$\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ $=\sqrt{2}{U^{\prime}}\left(-\frac{p_{x}p_{y}}{p_{0}^{2}}\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}+\frac{p_{z}}{p_{0}^{2}}\Im\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}\right)$ (84)
@@
With zero helicity injection, and if all homochiral-wave interactions are resonant ($F_{k_{f}}=1$), the total energy balance between the small-scale flux $\Pi_{\boldsymbol{p}}$ (83) and the transfer to the mean flow from layers II is therefore:
$\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{p_{x},p_{z}}^{\rm II}\equiv\int dp_{y}\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{\rm II}$ $=\epsilon_{p_{x},p_{z}}-\Pi_{p_{x},K_{y},p_{z}},$ (86)
@@
We now extract the leading order functional form of the Reynolds stress in the variable $S=\frac{{U^{\prime}}}{\Omega}$. Here we consider that $S$ is sufficiently large ($S\gtrsim l_{f}/L_{y}$, following the computation in [26]) so that all the waves excited at the forcing scale are coupled with the condensate via homochiral interactions, hence $F_{k_{f}}=1$. We also assume that
@@
The energy transfer between the waves and the 2D condensate in (105) decreases with decreasing rotation, via increasing $S={U^{\prime}}/\Omega$, due to the growing number of modes in sector A that break conservation of $H_{\boldsymbol{p}}^{\pm}$ and extract mean-flow energy. This energy extraction is larger as $L_{x}/l_{f}$ increases, due to vortex-tilting preferentially energizing modes with large $p_{z}/p_{x}\sim L_{x}/l_{f}$, via the term in $p_{z}/p_{x}$ in (87).
@@ E.2.1 Reynolds stress in Layers I
$\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ $=\frac{\epsilon_{p_{x},p_{z}}}{2\sqrt{2}p_{x}{U^{\prime}}}\frac{p_{0}^{2}k_{f}^{2}}{p^{4}},$
$\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}$ $=i\frac{\epsilon_{p_{x},p_{z}}}{2\sqrt{2}p_{x}{U^{\prime}}}p_{0}\frac{p_{z}}{p_{x}}\frac{k_{f}^{2}}{p^{2}}\Bigg[\arctan\Big(\frac{p_{y}}{p_{0}}\Big)+\arctan\Big(\frac{p_{y}^{f}}{p_{0}}\Big)\Bigg],~~~~~~~-p_{y}^{f}<p_{y}<-p_{y}^{*}$ (108)
@@
$\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s}=\frac{\epsilon_{p_{x},p_{z}}}{2p_{x}{U^{\prime}}}\frac{k_{f}^{2}}{p^{*}p},$ (109)
using the continuity of the flux at $p_{y}^{*}$, $\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s}=\frac{p^{*2}}{p_{0}^{2}}\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}(-p_{y}^{*})$. Because Eq. (53) exhibits $p_{y}\to-p_{y}$ symmetry, the boundary condition at $p_{y}^{*}$ dictates the values of the correlators at $p_{y}^{*}$.
@@
$\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}=\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}(-p_{y}^{*})\frac{p^{*4}}{p^{4}}+\frac{\sqrt{2}\epsilon}{2|{U^{\prime}}p_{x}|p^{4}}\int_{p_{y}^{*}}^{p_{y}}q^{4}(1-\frac{q_{y}^{2}}{q^{2}})\chi_{\boldsymbol{q}}dq_{y}=\left\{\begin{aligned} &\frac{\epsilon_{p_{x},p_{z}}}{2\sqrt{2}|{U^{\prime}}p_{x}|}\frac{p_{0}^{2}k_{f}^{2}}{p^{4}},~~~p_{y}^{*}<p_{y}<p_{y}^{f}\\ &\frac{\epsilon_{p_{x},p_{z}}}{\sqrt{2}|{U^{\prime}}p_{x}|}\frac{p_{0}^{2}k_{f}^{2}}{p^{4}},~~~p_{y}>p_{y}^{f}\end{aligned}\right$. (110)
$\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}=\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}(-p_{y}^{*})\frac{p^{*2}}{p^{2}}+\frac{i}{p^{2}}\int_{p_{y}^{*}}^{p_{y}}\frac{p_{z}}{p_{x}}q^{2}v_{q}^{2}dq_{y}=\left\{\begin{aligned} &i\frac{\epsilon_{p_{x},p_{z}}~p_{0}}{2\sqrt{2}{U^{\prime}}p_{x}}\frac{p_{z}}{p_{x}}\frac{k_{f}^{2}}{p^{2}}\Bigg[\arctan(\frac{p_{y}}{p_{0}})+\arctan(\frac{p_{y}^{f}}{p_{0}})-2\arctan(\frac{p_{y}^{*}}{p_{0}})\Bigg],\\ &~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~p_{y}^{*}<p_{y}<p_{y}^{f}\\ &i\frac{\epsilon_{p_{x},p_{z}}p_{0}}{\sqrt{2}{U^{\prime}}p_{x}p_{y}^{f}}\frac{p_{z}}{p_{x}}\frac{k_{f}^{2}}{p^{2}}\Bigg[\arctan(\frac{p_{y}}{p_{0}})-\arctan(\frac{p_{y}^{*}}{p_{0}})\Bigg],\\ &~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~p_{y}>p_{y}^{f}\end{aligned}\right$.
@@
$\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}=\left\{\begin{aligned} &\frac{\epsilon_{p_{x},p_{z}}k_{f}^{2}}{\sqrt{2}U^{\prime}}\Bigg[\frac{p_{y}}{p^{4}}-\frac{p_{z}^{2}}{p_{x}^{2}}\frac{1}{p_{0}}\frac{1}{p^{2}}\Big(\arctan\Big(\frac{p_{y}}{p_{0}}\Big)+\arctan\Big(\frac{p_{y}^{f}}{p_{0}}\Big)\Big)\Bigg],~~~~~-p_{y}^{f}<p_{y}<-p_{y}^{*}\\ &\frac{\epsilon_{p_{x},p_{z}}k_{f}^{2}}{\sqrt{2}U^{\prime}}\Bigg[\frac{p_{y}}{p^{4}}-\frac{p_{z}^{2}}{p_{x}^{2}}\frac{1}{p_{0}}\frac{1}{p^{2}}\Big(\arctan(\frac{p_{y}}{p_{0}})+\arctan(\frac{p_{y}^{f}}{p_{0}})-2\arctan(\frac{p_{y}^{*}}{p_{0}}\Big)\Bigg],~~~p_{y}^{*}<p_{y}<p_{y}^{f}\\ &\frac{\sqrt{2}\epsilon_{p_{x},p_{z}}k_{f}^{2}}{U^{\prime}}\Bigg[\frac{p_{y}}{p^{4}}-\frac{p_{z}^{2}}{p_{x}^{2}}\frac{1}{p_{0}}\frac{1}{p^{2}}\Big(\arctan(\frac{p_{y}}{p_{0}})-\arctan(\frac{p_{y}^{*}}{p_{0}}\Big)\Bigg],~~~~~~~p_{y}>p_{y}^{f}\end{aligned}\right$. (111)
In Fig. 1e in the main text, we visualize ${U^{\prime}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ in the $(p_{y},p_{z})$ plane by showing Eq. (111) in layers I, completing the Reynolds stress in layers II ( Eq. (85)). In the figure, we fix $p_{x}=8\pi/L_{x}$ and $S=1$.